OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line through the midpoint of and perpendicular to the segment joining points (1, 0) and (5, -2).

OpenStudy (anonymous):

@apoorvk

OpenStudy (anonymous):

i know the mid-point is (3 , -1).

OpenStudy (apoorvk):

yeah great, now can you find out the equation of the line passing through those two points?

OpenStudy (anonymous):

y+1=1/2(x-3) is that right?

OpenStudy (anonymous):

or wrong equation?

OpenStudy (anonymous):

well the slope of the line perpendicular to d line joinin d points is 2 nd nt 1/2

OpenStudy (anonymous):

sorry. what?

OpenStudy (anonymous):

the equation is (y+1)=2(x-3)

OpenStudy (anonymous):

but i thought the slope was -1/2... y2-y1/x2-x1 so -2-0/5-1 is -2/4 which is -1/2

OpenStudy (anonymous):

that is the slope of d line joinin d points.any line perpendicular 2 dis line will have a slope which is d negative reciprocal of d earlier slope! the product of d slopes of perpendicular lines are always = -1

OpenStudy (anonymous):

okay. (sorry i just get confused when people don't speak english.) so the slope of the perpendicular line is 2?

OpenStudy (apoorvk):

Ohkay, what you did first was right, Mr. Sailor!

OpenStudy (apoorvk):

The equation of the line passing through the two points is exactly what you found out. Now, express that in standard form, i.e. 'ax + by + c = 0', can you?

OpenStudy (anonymous):

y+1=1/2(x-3) this? or (y+1)=2(x-3)

OpenStudy (apoorvk):

The first one. That is: \[y + 1 = \frac 1 2 (x-3)\]

OpenStudy (anonymous):

2y + 2 = 2x - 6 2y = 2x - 8 y = x - 4 is that right so far?

OpenStudy (anonymous):

yes, no?

OpenStudy (apoorvk):

Umm, you sent that '2' in the denominator to the other side, okay, but then you have 2y + 2 = x- 3 right? :D

OpenStudy (anonymous):

ah. yes. okay. i see. so then 2y + 2 = x - 3 2y = x - 5 -x + 2y= -5 i feel i went wrong somewhere. :/

OpenStudy (anonymous):

x - 2y = -5?

OpenStudy (apoorvk):

yeah great now! Now now.. For a straight line "ax + by = c", the equation of the line perpendicular to it is: "bx - ay = k" where 'k' is another constant. So, what do you think would be the equation of the perpendicular line in terms of 'x', 'y' and 'k'?

OpenStudy (apoorvk):

No, sailor, the r[previous one was correct that is '- x + 2y = -5'

OpenStudy (anonymous):

wait wait wait. where did the a and the b come from?

OpenStudy (apoorvk):

It's the standard equation of a straight line!

OpenStudy (apoorvk):

compare and find out 'a' and 'b' in in your case.

OpenStudy (anonymous):

oh my brain. okay. -1 is a and 2 is b?

OpenStudy (anonymous):

and k is -5?

OpenStudy (apoorvk):

great! no the constants are different, 'k' and is not the same as 'c'. so, what would be the equation of the perpendicular line if it's "bx-ay =k" forget 'k' for a moment.

OpenStudy (apoorvk):

I mean include 'k' in the equation but write it just as 'k'.

OpenStudy (anonymous):

2x+y=k

OpenStudy (apoorvk):

great! now this line passes through that midpoint (3,-1). so that means this point satisfies the equation of the perpendicular line. So substitute and find out 'k' from this, can you?

OpenStudy (anonymous):

2*3 + 1 =k k=7

OpenStudy (apoorvk):

yeahh!! so you got 'k' now substitute it in the equation for the perpendicular line that we had obtained, i.e.: 2x+y=k that would be your answer!

OpenStudy (anonymous):

2x+y=7 thanks! just wondering, is there a faster way of doing this?

OpenStudy (apoorvk):

yeah, do this without my help, and practice it a bit, this is basically just a two step process if you notice.

OpenStudy (anonymous):

alright:) thanks again

OpenStudy (apoorvk):

No worries ;)